OpenStudy (anonymous):

@perl Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

OpenStudy (anonymous):

@perl There is three questions on this test that i am stuck on and this is the 2nd one

OpenStudy (anonymous):

btw its 

OpenStudy (anonymous):

@itsbribro @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

@dtuniyants

OpenStudy (itsbribro):

@Catlover5925

OpenStudy (perl):

try one with d = 0, and pick c so that it is negative

OpenStudy (anonymous):

so it would be a square root x+b+(-1)=0?

OpenStudy (perl):

my idea was 1*sqrt(x + 2) -1 = -1 if you add 1 to both sides you get 1*sqrt( x+2) = 0 that has only one solution if you square both sides (x+2) = 0 x = -2

OpenStudy (perl):

this equation does not have an extraneous solution

OpenStudy (anonymous):

dang your fast.

OpenStudy (anonymous):

it shows show why it is extraneous

OpenStudy (perl):

actually any positive value will work

OpenStudy (perl):

1*sqrt(x + 2) -1 = 3 works as well, there is no extraneous solution but if you do 1*sqrt(x+2) -1 = -3 , you get an extraneous solution. do you see why?

OpenStudy (perl):

try to solve 1*sqrt(x+2) - 1 = -3 and then plug in the solution

OpenStudy (anonymous):

after i add one do i square root both sides?

OpenStudy (perl):

you can square both sides

OpenStudy (anonymous):

when i square root -2 i get an error message

OpenStudy (perl):

you should square both sides , not square root

OpenStudy (perl):

1*sqrt(x+2) - 1 = -3 sqrt(x+2) = -2 we want to solve for x, so we square both sides sqrt( x+2)^2 = (-2)^2 x+ 2= 4

OpenStudy (anonymous):

oh okay i got it now

OpenStudy (perl):

now if you plug in that solution back into the original equation, what happens?

OpenStudy (anonymous):

hold on what is an extranous solution?

OpenStudy (anonymous):

It equals itself right? @perl

OpenStudy (perl):

nope

OpenStudy (anonymous):

x would equal 2 right?

OpenStudy (perl):

an extraneous solution is like an illegal solution , it comes extra when you solve for x

OpenStudy (perl):

lets plug in x = 2 into the original equation 1*sqrt(x+2) - 1 = -3 1*sqrt( 2 + 2 ) -2 = sqrt(4) -2 = 2 - 2 = 0 but 0 does not equal to -3

OpenStudy (anonymous):

okay

OpenStudy (perl):

so it turns out x = 2 is an invalid solution, so we call it extraneous . we got it when we tried to solve for x by squaring

OpenStudy (perl):

in general, when you square an equation, you can introduce an invalid or extraneous solution

OpenStudy (anonymous):

so that is what i would put for part 3 and then my work for part 2?

OpenStudy (anonymous):

so for the first one that is not extranous what happens when you plug in -2?

OpenStudy (perl):

but the equation 1*sqrt(x + 2) -1 = 3 this does not have an extraneous solution. when we solve it we get sqrt(x+2) = 4 (sqrt(x+2))^2 = 4^2 (x+2) = 16 x = 14 if you plug in x = 14 into original equation, it satisfies it

OpenStudy (anonymous):

you said the equation was 1*sqrt(x+2)-1=-1

OpenStudy (perl):

yes i kind of changed it :0

OpenStudy (perl):

that equation works as well if you plug in x = -2 into the equation 1*sqrt(x +2) -1 = -1 you get 1*sqrt( -2 + 2 ) -1 = sqrt(0) -1 = 0 -1 = -1 so it works and x=-2 is not an extraneous solution

OpenStudy (anonymous):

okay i got one last question that i tagged you in!